A paper cup is a perfect cylinder that is 6.1 cm tall and 4 cm across. Find the total surface area of the cup to the nearest square into. Use 3.14 for pi.
Answer by hlkitty96(1)

S=2πrh + 2πr^2
R=1/2d
R=1/2(4)
R=2
S=2π(2)(4.1) + 2π(2)^2
S=2π(12.2) + 2π(4)
S=24.4π + 8π
S=76.62 + 25.12
S=101.74 cm^2
I rounded to the nearest tenth.